APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Tuesday May 29, 2007 (MJD 54250) ---=== OBSERVING TEAM ===--- Audrey Simmons Kaike Pan Brian Lee (UF) Wendell Jordan (support by phone) ---=== OBSERVING PLAN ===--- ET Science! ---=== OBSERVING SUMMARY ===--- The skies were clear and photometric all night. We had trouble accessing the cloud camera images on the webpage. The graphs were fine, but the ircam image came up as a question mark. Can someone take a look at this to see what is wrong? During afternoon checkout, we had trouble clearing the axes in the TCC and MCP windows because of a interlock bypass problem. Wendell was called. Hitting a reset buttom solved the problem (see log for more details). We had difficuity to acquire field for cartridge 7 (plate 8022). 50 minutes was billed to it, and we ended with giving up on this cartridge altogether (see log for more details). However, we were able to get good exposures on: Cart 3, plate 8024 Cart 5, plate 8025 (2 exposures) Cart 1, plate 8026 Cart 9, plate 8028 Cart 6, plate 8027 Cart 2, plate 8032 ---=== OBSERVING LOG ===--- Afternoon: ---------- We were not able to clear axes with "axis init" commands from tcc session or clear command from mcp window during afternoon checkout. When one axis cleared, the "AB" for azimuth and altitude axes came back right way. Interlocks showed the brakes for these two axes were on. Wendell was called to ask of something has been done today or knew about what was going on. He said it seemed to be an interlock problem, but we checked the interlock bypass panel dowstairs about 20 min prior to calling. There was no axis slip or anything abnormal and the interlock message did not give us more information about why the azimuth and altitude brakes were on. We went downstairs and confirmed that all lights in interlock bypass panel were still normal and that no axis slips were detected. We could not think anything else, we decided to hit the reset button anyways to reset all axes. Then, our next attempt to clear the axes was successful. 01:10Z - Started ventilation with doors are louvers open, fans on. Night: ------ 02:05Z - Opened enclosure. 03:07Z - Cart 3, plate 8024 mounted. GSFOGTF. 50 min exposure taken. Seeing 1.5". After exposure finished, a I2 template was taken with calibration lamps. 04:18Z - Cart 5, plate 8025 mounted. GSFOGTF. One 50 min exposure and one 20 min exposure taken. Seeing variable from 1.9-2.3". 05:33Z - Cart 7, plate 8022 mounted. We found no guide stars on goToField and no fk5 star in fiber 10/11 (including applied offsets in A and D). We checked fiducial crosses and confirmed that they were good. We attempted to use the offsets that were used a few nights ago for this cartridge, but still no stars were found. Then we cleared the offsets, and didnt have any luck with fk5InFiber, one spiral pattern 5 10 2, another spiral pattern 10 10 3 either (however, we did find a star but it was to faint to be an fk5 star). We tried the whole process over again by clearing offsets, and applied the offsets from a few nights ago, ran another spiralPattern 10 10 3, but still did not find the fk5. Note that the field has the center coordinates: ra -- 247.6, dec -- 38.4 and we were acquiring the field about 1.5 hour before its transit point. We thought it should be okay to acquire the field at the moment, and the plate was observed two nights ago. We noticed that the rotator was moving at a rate about twice fast as the other cartridge (cart 1 - our next plate), the azimuth was moving also faster than other plates (2 times fast as cartridge 1 plate). This plate has been observed twice in these ET run and neither of these observations needed an fk5 to acquire the field. Is it possible the coordinates for fk5 in the mapping files were not accurate? After 50 min, we gave up this cartridge, and moved on with cartridge 1. 50 minutes was billed to the difficulty to acquire the field. 06:26Z - On cartridge 1 (plate 8026), we found the stars quicker than the last cartridge. However, we had to use a spiral pattern to find the offsets (with a clearing offsets): 10" in A, and -10" in D. So this means that the telescope is probably not the problem. Started 50 min exposure. Seeing 1.5-2..2". 07:34Z - ET mid-calibrations. 07:50Z - Cart 9, plate 8028. GSFOGTF. 50 min exposure. Seeing 1.3". 08:55Z - Cart 6, plate 8027. GSFOGTF. 50 min exposure. Seeing 1.5". 10:00Z - Cart 2, plate 8032. fk5 star was used to find field. 50 min exposure. Seeing 1.5". 11:00Z - Enclosure on, doors closed, fans off. 11:09Z - endNight rejoiced ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): QA Procedures Done (y/n): UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 45092, plate -9999 ------- 00:22 45092 0.0 bias ----- sequence 45093, plate 8032 ------- 00:54 45093 10.0 flat 00:56 45094 2.0 arc ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 03:28Z 3 8024 281.36 -0.0001 50.52 0.0049 119.77 -0.0111 1.000260 04:26Z 5 8025 -51.26 -0.0018 47.94 0.0033 138.89 -0.0346 1.000150 06:49Z 1 8026 221.53 0.0008 54.71 0.0052 70.10 -0.0054 1.000260 08:30Z 9 8028 140.05 0.0023 72.53 0.0059 -50.12 -0.0016 1.000110 08:56Z 6 8027 157.50 0.0072 78.27 0.0044 -26.40 0.0154 1.000140 10:02Z 2 8032 77.76 0.0021 52.99 0.0070 240.26 0.0047 1.000170 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 03:28Z 3 8024 1.00026 -2568 -1072 -125 281 50.5 12.3 21 236 - 1.5 04:26Z 5 8025 1.00015 -1481 -189 -150 -51 47.9 12.6 14 248 - 2.0 06:49Z 1 8026 1.00026 -2568 -984 -25 222 54.7 12.8 21 294 - 1.5 08:30Z 9 8028 1.00011 -1086 203 -25 140 72.5 12.1 10 314 - 1.3 08:56Z 6 8027 1.00014 -1383 -49 -25 157 78.3 12.0 0 57 - 1.5 10:02Z 2 8032 1.00017 -1679 -247 -30 78 53.0 12.0 5 358 - 1.5 ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 23:14Z 64 23 20 255 (WSW) 791 - Clear 23:46Z 61 20 22 248 (WSW) 781 - " 00:19Z 62 20 21 258 (WSW) 711 - " 00:51Z 60 15 16 272 (W) 753 - " 01:21Z 59 15 15 263 (W) 750 - " 01:51Z 57 15 14 244 (WSW) 779 - " 02:21Z 55 15 16 235 (SW) 746 - " 02:53Z 54 14 19 234 (SW) 761 - " 03:23Z 54 14 18 253 (WSW) 701 - " 03:55Z 54 13 14 252 (WSW) 763 - " 04:26Z 54 13 13 247 (WSW) 686 - " 04:59Z 55 12 12 265 (W) 657 - " 05:30Z 55 11 17 288 (WNW) 545 - " 06:02Z 55 11 18 272 (W) 609 - " 06:33Z 54 8 12 276 (W) 584 - " 06:57Z 54 10 16 293 (WNW) 622 - " 07:29Z 54 11 17 296 (WNW) 646 - " 08:00Z 54 11 21 302 (WNW) 640 - " 08:30Z 53 12 10 313 (NW) 605 - " 09:07Z 53 12 7 336 (NNW) 516 - " 09:40Z 53 11 7 344 (NNW) 532 - " 10:12Z 53 11 6 350 (NNW) 529 - " 10:44Z 52 11 5 358 (NNW) 515 - " ---=== TELESCOPE STATUS ===--- 01:10Z - Began ventilation. Doors and louvers opened, fans on. 02:05Z - Opened enclosure. 11:00Z - Enclosure is on, doors and louvers are closed, and fans are off. No interlocks bypassed. Status at: 11:08Z Telescope stowed at: 30,121,00 mount Instrument mounted: Cartridge 2 Counterweights at: 250 lbs LN2 autofill systems: Connected and turned on. 180L LN2 dewar scales: Spectro 141.60 lbs, 21.85 psi Imager 113.50 lbs, 17.06 psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_34_0 Watcher: v2_36_0 MCP: v6_4_0 TPM: tpm_v3_5_0 AstroDa: v15_13_1 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: v2_30 SoS: v5_1_8 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A 0.0080 -0.0050 Axial B 0.0060 0.0090 Axial C 0.0090 0.0160 Trans D 0.0080 0.0040 Lateral E 9.9695 9.9568 Lateral F 11.5443 11.5443 GALILS Commanded: 4600. -2000. 2650. -9300. -9800. -9800. Actual: 4600. -2000. 2650. -9300. -9800. -9800. SETMIR VALUES PriDesOrient: 0.00 0.00 0.00 0.00 0.00 PriOrient: 0.00 0.00 0.00 0.00 0.00 SECONDARY: ---------- Focus: 0.00 Air Temp.: 13.1 Alt.: 29.996450 MIGS TONIGHT NOMINAL Axial A -0.0300 -0.0240 Axial B -0.0280 -0.0280 Axial C -0.0430 -0.0370 Trans D -0.0100 -0.0040 GALILS Commanded: 1614055. 1594748. 1543158. -10550. -5800. Actual: 1614053. 1594719. 1543082. -10535. -5822. SETMIR VALUES SecDesOrient: 1257.00 20.00 -10.00 -100.00 176.64 SecOrient: 1257.03 19.98 -9.99 -100.43 176.57 ---=== PROBLEMS IN DETAIL ===--- Trouble with accessing the ircamera images on the webpage. Interlock bypass problem causes us to have trouble clearing the axes in the TCC/MCP windows during afternoon checkout.